# General Place Recognition Survey: Towards Real-World Autonomy

### v1

@@ Abstract
This paper begins with an exploration of PR’s formulation and key research challenges. We extensively review literature, focusing on related methods on place representation and solutions to various PR challenges. Applications showcasing PR’s potential in robotics, key PR datasets, and open-source libraries are discussed. We also emphasizes our open-source package, aimed at new development and benchmark for general PR. We conclude with a discussion on PR’s future directions, accompanied by a summary of the literature covered and access to our open-source library, available to the robotics community at: https://github.com/MetaSLAM/GPRS.
@@ I Introduction
Figure 1: Structure of our GPR Survey. Place Recognition (PR) is the ability to recognize visited areas under different environmental conditions and viewpoint differences. In this survey, Section II defines the problem of position-based PR and introduces the significant challenges. Section III investigates methods in place representation. Section IV and Section V provide the solutions for the current four major challenges and potential applications, respectively, Finally, Section VI introduces the current datasets, metrics, and related supported libraries for PR research.
Figure 2: This timeline the evolution of Place Recognition (PR), transitioning from the handcrafted methods to data-driven methods. It also highlights key features of representative PR methods, summarizes related surveys in this field, and introduces applications along with the development of PR approaches. Our survey emerges at an opportune moment, given mojor events across multiple fields including AI, reconstruction, and collborative perception.
@@ I-A Background
As the foundamental module in navigation, SLAM enables a robot to estimate its ego-motion while simultanously constructing a map of its surroundion environments. As a distant yet exciting frontier in the community, lifelong navigation raises new performance requirements of SLAM beyond accuracy from several aspects including large-scale and long-term localization, map maintainance in dynamic environments, active perception, and self-evolution. This indicates that we are entering the SLAM2.0 era: distinguished from traditional SLAM [1], SLAM2.0 emphasizes the integration of advanced algorithms in robotics and AI to support the scalable, adaptable, and efficient realization of lifelong robotic autonomy.
As the core of advancing to SLAM2.0, PR, now is becoming more essential than ever, which enables robots to identify previously visited areas despite changes in environmental conditions and viewpoints. Essentially, PR’s ability hinges on creating short-term or long-term association between current observations and a robot’s internal “memory”. In visual SLAM (VSLAM), “memory” typically refers to a database of stored visual information. For decision making, PR enables robots to apply lessons from past experiences to current situation. Thus, PR’s role extends beyond merely loop closure detection (LCD) in SLAM. Its applications now span failure recovery, global localization, multi-agent coordination, and more.
Over the years, PR methodologies have evolved significantly, as evidenced by facts such as: (1) there are more than $\mathbf{3400}$ published papers within this area; (2) the complexity of the operational scope for robots escalates markedly when transitioning from controlled indoor environments to the unpredictable conditions of outdoor fields, which brings new performance requirements; and (3) there have been several massive PR competitions, such as the CVPR 2020 Long-term Visual PR (VPR) competition and our previous ICRA 2022 General PR competition for City-scale UGV Localization and Visual Terrain Relative Navigation (VTRN). Thus, PR has drawn huge interest from the robotics and computer vision (CV) research communities, leading to a large number of PR techniques proposed over the past many years.
@@ I-B Summary of Previous Surveys
The evolution of PR visualized as the timeline in Fig. 2. A comprehensive historical analysis up to the year $2015$ on the PR issue is meticulously provided by Lowry et al.[2], emphasizing vision-based methodologies. This seminal survey articulates the VPR challenge, delineates the core components of recognition systems, and reviews solutions to tackle variations in appearance. It distills the essence of a PR system into three fundamental modules: (1) an image processing module for visual input abstraction, (2) a map representing the robot’s understanding of the world, and (3) a belief generation module to evaluate the likelihood of the robot being in a previously encountered or new location. This architecture has served as the foundations of modern PR solutions. The evolution in PR methodology has been profound, transitioning from manual feature extraction to deep learning-based approaches, as comprehensively discussed in prominent studies [3, 4, 5]. However, as Zaffar et al. [6] note, the PR field has seen a growing divergence, particularly as the robotics and CV communities have developed distinct performance benchmarks for PR. This divergence has rendered direct comparisons between PR methods problematic, owing to the significant variability in benchmark datasets and evaluation metrics. To address this issue, they introduced an open-source, standardized framework for PR evaluation, with a particular emphasis on VPR [6]. And there are related works concentrate on visual [7] or LiDAR localization [8], where several literation on VPR and LiDAR-based PR (LPR) are covered.
@@ I-C Contributions and Paper Organization
Given new approaches and applications along with development of robots and AI, the field of PR is active. However, existing literature lacks comprehensive surveys that explore the diverse aspects of place representations, the underlying challenges, and the applications that highlight potential opportunities in place recognition research. In contrast, our goal is to address and fill this gap. Additionally, this paper emphasizes “General PR” (GPR) as it extends beyond VPR and LPR to cover a broader range of topics within the field. This includes the utilization of multi-modal information, such as language, graphs, and implicit embeddings, to enhance PR. The timing for such a survey is particularly fortuitous given major events across these related fields: for example, the emerging large-language models (LLM) such as the ChatGPT [9] exhibit a remarkable ability, neural mapping-related methods [10] evolve the way of environmental representation and scene rendering, and widespread presence in robots. In overall, this paper is structured in Fig. 1 and summarized as follow:
  • Section II details two widely accepted definitions of PR: the position-based and the overlap-based approaches. Adhering to the position-based definition to narrow the focus, this paper then proceeds to offer a more precise problem definition (Section II-A) and to highlight the key challenges involved (Section II-B).
  • Section III reviews existing representation approaches in PR, covering the core solutions prevalent in the field. Intuitively, PR extends beyond mere image-based approches, encompassing a variety of solutions. At a low level, detailed in Section III-A, a “place” can be captured through sensors, such as cameras, LiDARs and Radars. It’s generally expected that identical locations will produce similar sensor data. On a more abstract and higher level, as discussed in Section III-B, a “place” may also be represented through compact data forms, like scene graphs and implicit embeddings.
@@
  • Section VI reviews the leading datasets and benchmarks in the field of PR. It introduces a new perspective on property analysis that complements to primary metrics for quality assessment. Addtionally, this section showcases open-source PR libraries, featuring our in-development library dedicated to GPR research.
@@
II Definition of Place Recognition and Challenges
@@
Before exploring specific solutions in PR, it is crucial to address two basic questions:
  1. 1.
    What is the rigorous definition of PR?
  2. 2.
    What are the primary challenges encountered in PR?
@@
Figure 3: In position-based PR, the focus is on identifying if a query image captures the same location as a reference image. For instance, in the provided images, position-based PR would recognize the first query image as matching the reference place, but deem the second query that is taken from a geographically distant location as a different place. Overlap-based PR, however, would classify both query images as the same place as they share visual overlap with the reference, denoted by the red box. But the second query, captured from a distant position, offers limited utility for navigation.
II-A Definition: What is Place Recognition?
There primarily exist two prevalent PR definitions: position-based and overlap-based. They diverge in their definitions of a “place”, evaluation metrics, objectives. Fig. 3 visualizes the key differences between these definitions.
O’Keefe’s identification of ”place cells” within the hippocampus, a finding that contributed to winning the Nobel Prize in 2014 [11] in neuroscience, inspiring the concept of PR in robotics. Building on this discovery, Lowry et al. defined the position-based PR as a robot’s capability to identify previously visited areas, despite changes in environmental conditions or viewpoint [2]. This perspective treats a “place” akin to a region, though its precise definition can vary with the context of navigation-sometimes seen as a specific location, other times as a broader area. Essentially, the challenge of PR revolves around determining if the current location is within proximity to a previously encountered region. Alternatively, Garg et al. introduce the concept of overlap-based PR, where the focus is on identifying potential matches through visual similarities [12]. According to this approach, even if an exact location is viewed from markedly different angles, it’s considered a distinct place. This method closely parallels the task of content-based image retrieval in the computer vision field [13], which involves searching a large database to find the image most similar to a provided query image.
The overlap-based approach aligns with the visual retrieval methods prevalent in the computer vision community. However, from a robotics standpoint, these methods don’t seamlessly translate to SLAM and navigation due to the unclear relationship between visual overlap and the precise position of retrievals. This ambiguity is exemplified in Fig. 3, where despite the top-left and bottom images viewing the same landmark (i.e., the building) within their overlapping field of view (FoV), determining their relative transformation based on visual overlap alone is challenging. For subsequent tasks such as pose graph optimization (PGO) [14] that focuses on maximizing inlier ratios, the position-based definition is clearer and more advantageous. Consequently, this paper will focus on this definition, starting with its mathematical formulation and exploring relevant research in the field.
II-A1 Problem Formulation
A fundamental assumption underlying position-based PR is that two geographically close places share similar layout which can be partially observed by sensors such as cameras. Let $Q$ represent the current place of the robot, let $\mathcal{P}$ denote a set of previously visited places, and let $P\in\mathcal{P}$ indicate the geographically close place to $Q$. If a place’s layout can be abstracted as a global descriptor $\mathbf{g}$, the above assumption is formulated as
$\text{if}\ d(P,Q)<\delta\ \ \text{then}\ \ \mathbf{g}_{P}\approx\mathbf{g}_{Q},$ (1)
where $d(\cdot,\cdot)$ represents geometric distance like Euclidean distance between the place $P$ as well as $Q$ and $\delta$ is the threshold. The value of $\delta$ is manually set since the definition of a specific place depends on subsequent tasks. Modern position-based PR solutions, predominantly framed within end-to-end learning paradigm, follow the above assumption and optimize modal parameters for the most correct place $P^{*}$ such as
$P^{*}=\underset{P\in\mathcal{P}}{\arg\min}\ d(P,Q)\ \Rightarrow\ P^{*}=\underset{P\in\mathcal{P}}{\arg\min}\ \|\mathbf{g}_{P}-\mathbf{g}_{Q}\|,$ (2)
[…]
@@ II-B Challenges
    Viewpoint Difference: This issue arises from differences in how sensors capture an environment’s perspective. For example, a building’s frontal view exposes its full shape, whereas a top-down perspective highlights the layout and roof design. Such viewpoint variations are especially noticeable when a robot revisits a location from a different angle or altitude. This challenge is ubiquitous across sensor types, necessitating PR systems to incorporate robust modules for feature extraction and matching to accommodate these perspective shifts.
@@
    Generalization Ability: For lifelong navigation, the boundless complexity of environments makes generalization to unseen areas a critical challenge for PR methods. Additionally, these methods must support online learning to adapt over time, a crucial requirement for missions like space exploration, where robots investigate other planets long-term with minimal human supervision. Robots should possess the capability to continuously learn and adapt to new environments.
@@
    Efficiency on Resource-Constrained Platforms: Adapting PR algorithms that mainly use neural networks for robots with limited computational resources poses a significant challenge, particularly for drones. Effective algorithms should balance accuracy with computational efficiency. This becomes a necessity for both single and multi-robot systems which face bandwidth constraints and potential communication disruptions.
@@
Building from our overview of PR’s definitions and challenges, the next sections will delve into associate studies.
III Describing Places: Data Representation
@@
The pivotal aspect, and indeed the foundation of PR, lies in representing a place with compact features while preserving key distinctions. Therefore, this section delves into the various formats of representations employed in prevalent PR methods, ranging from low-level, sensor-specific representations to high-level, sensor-independent representations.
III-A Low-Level Representations
@@ III-A1 Sensor Selection Criteria
Cameras, LiDAR (Light Detection and Ranging), and Radar (Radio Detection and Ranging) are essential sensors in place recognition (PR). Important selection criteria include field of view (FoV), information density, and robustness under various conditions. Frame cameras provide high-resolution images, while event cameras perform better in low-light and reduce motion blur. LiDARs generate precise 3D point clouds but have lower resolution. Radars offer long-range capabilities and excel in poor weather, also measuring relative velocity via the Doppler effect. Selecting between cameras, LiDARs, and Radars depends on the specific needs for precision, range, and environmental suitability.
@@ III-A2 Camera-Related Approaches
VPR is the most thoroughly investigated problem. As highlighted in previous surveys (Section I-B), it has been comprehensively explored through both traditional and learning-based methodologies.
Handcrafted VPR Methods: Traditional handcrafted methods like[23], [24] are well reviewed and extensively used in commercial robotics systems. However, data-driven approaches, leveraging the power of deep-learning algorithms, consistently outperforms their counterparts in terms of accuracy and robustness. This superior performance has catalyzed data-driven approaches as the preferred choise for PR.
Data-Driven VPR Methods: Data-driven approaches, particularly those utilizing deep neural networks (DNNs), provide the unique benefit of automatically learning features directly from training data, shifting from the handcrafted methods that require significant engineering and domain expertise. The capability of DNNs to derive distinct features has been significantly improved, thereby advancing PR across various complex scenarios. Moreover, utilizing DNNs allows for the end-to-end training of PR models. The trend indicates that VPR tasks continue to adopt novel image learning methods.
The emergence of NetVLAD[25] marked a significant advance in data-driven methods, utilizing CNNs as local feature extractors and introducing differentiable VLAD (Vector of Locally Aggregated Descriptors) layers for feature aggregation. Subsequent developments enhanced both feature extractors and aggregators. Notably, Regional Maximum Activations of Convolutions (R-MAC)[26] and Generalized Mean (GeM) [27, 28] emerged as effective alternatives to VLAD for feature aggregation. With the rise of vision transformers, which surpass CNNs in performance, TransVPR[29] was developed, incorporating multi-level attention mechanisms to improve image understanding. However, the adoption of vision transformers significantly increases the model size, requiring more computational resources for end-to-end training. Foundation models, pre-trained on extensive datasets, demonstrate robust zero or few-shot capabilities in various vision tasks[30] and are directly applicable to VPR tasks. AnyLoc[31] utilizes the ViT-based DINOv2[30] to achieve state-of-the-art performance, underscoring the potential of foundation models.
@@
Various research efforts have broadened scope by integrating diverse cues such as semantic, geometric, event data to boost the place representation. The utilization of semantics includes the way of filtering specific pixels [32] and adjusting the weight of feature embeddings [33]. Geometric cues, such as the 3D position of landmarks and ego-motion obtained from monocular image sequences or multiple cameras, also contribute to PR accuracy [34, 35]. Oertel et al. [34] introduced a fusion-based VPR method that leverages both 2D imagery and 3D points from structure-from-motion, Furthermore, event cameras, surpassing frame cameras in dynamic range, prove advantageous for VPR in fluctuating lighting conditions. Lee et al. [36] proposed to use event cameras to capture texture information under low-light condition, constructed edge-based images from event data to achieve PR.

### v2

@@ Abstract
This paper begins with an exploration of PR’s formulation and key research challenges. We extensively review literature, focusing on related methods on place representation and solutions to various PR challenges. Applications showcasing PR’s potential in robotics, key PR datasets, and open-source libraries are discussed. We conclude with a discussion on PR’s future directions and provide a summary of the literature covered at: https://github.com/MetaSLAM/GPRS.
@@ I Introduction
Figure 1: Structure of our General Place Recognition (PR) Survey. PR is the ability to recognize visited areas under different environmental conditions and viewpoint differences. This survey is structured as follows: Section II defines the problem of position-based PR and introduces the significant challenges. Section III investigates methods in place representation. Section IV and Section V provide the solutions for the current four major challenges and potential applications, respectively, Finally, Section VI introduces the current datasets, metrics, and related supported libraries for PR research.
Figure 2: This timeline maps the evolution of PR from handcrafted to data-driven methods, analyzing key techniques, surveys, and applications. Our survey emerges at an opportune moment, given major events across multiple fields including embodied AI, reconstruction, and collaborative perception. Citations for select works are omitted due to limited page; readers may search referenced titles online.
@@ I-A Background
As the fundamental module in navigation, SLAM enables a robot to estimate its ego-motion while simultaneously constructing a map of its environment. Lifelong navigation, the capability for robots to autonomously adapt to ever-changing environments with increasing experience, drives SLAM systems beyond accuracy-focused metrics, demanding solutions for long-term localization, dynamic mapping, and self-evolution. This signals the dawn of the SLAM2.02.02.02.0 era: diverging from traditional SLAM frameworks [1], SLAM2.02.02.02.0 prioritizes the synergy of advanced techniques such as machine learning-driven perception and semantic scene abstraction to enable scalable and efficient robotic autonomy in dynamic, open-world environments.
As the core of advancing to SLAM2.02.02.02.0, PR is now becoming more essential than ever, which enables robots to identify previously visited areas despite changes in environmental conditions and viewpoints. Essentially, PR’s ability hinges on creating short-term or long-term association between current observations and a robot’s internal “memory” of environments. In visual SLAM (VSLAM), memory typically refers to a map (or called database in some context) that consists of visual information. For decision-making, PR enables robots to associate and apply past experiences to the current situation. Thus, PR’s role extends beyond merely loop closure detection (LCD) in SLAM. Its applications now span failure recovery, global localization, multi-agent coordination, and more.
PR has emerged as a cornerstone capability for robotic autonomy, driven by three key indicators of its growing importance: (1) Over $\mathbf{3500}$ papers have been published on PR methodologies to date; (2) the transition from controlled indoor settings to unstructured outdoor environments, with challenges like perceptual aliasing and dynamic conditions, has necessitated robust, scalable PR solutions; and (3) the field has seen a surge in organized competitions and workshops, including the CVPR 2020 Long-term Visual PR (VPR) Challenge, the ICRA 2022 General PR competition for City-scale UGV Localization and Visual Terrain Relative Navigation (VTRN), and the IROS 2023 Closing the Loop on Localization Workshop. These converging factors position PR as a critical frontier in autonomous systems, catalyzing decades of methodological innovation spanning geometric, semantic, and learning-based paradigms.
@@ I-B Summary of Previous Surveys
The evolution of PR is visualized as the timeline in Fig. 2. A comprehensive historical analysis up to the year $2015$ on VPR is provided by Lowry et al. [2]. This seminal survey articulates the VPR challenge, delineates the core components of a PR system, and reviews major solutions to tackle appearance changes. It distills the essence of a PR system into three fundamental modules which still serves as the foundations of modern solutions: (1) an image processing module for visual input abstraction, (2) a map to represent how a robot stores its memory of the environment, and (3) a belief generation module to evaluate the likelihood of the robot being in a previously encountered or new location.
In the past decade, the evolution of mainstream PR methods has been transitioned from handcrafted descriptors to data-driven pipelines, as comprehensively discussed in prominent studies [3, 4, 5]. However, as Zaffar et al. [6] observed, the PR community has developed increasing fragmentation in performance benchmarking. This disparity complicates direct method comparisons due to inconsistent evaluation metrics and dataset implementations across studies. To resolve these challenges, their work established an open-source standardized evaluation framework specifically focused on VPR [6]. Other surveys like [7] (vision-based) and [8] (LiDAR-based) exclusively address VPR and LPR respectively. While these works review pose estimation techniques for fine-grained metric localization, such methods fall outside our core scope.
@@ I-C Contributions and Paper Organization
Recent progress in autonomous systems and machine perception has driven dynamic evolution in PR research. However, existing literature lacks a survey that thoroughly explores the diverse aspects, challenges, and deployment potentials for embodied AI applications. Our work addresses this gap by introducing a “General PR” (GPR) framework, which extends beyond the scope of VPR and LPR to encompass a wider spectrum of topics. GPR emphasizes the utilization of multi-model information such as visual, geometric, and textual inputs to establish robust environmental embeddings. This systematic review is particularly pertinent given transformative developments in three key areas: (1) Foundational language models [9, 10] demonstrating unprecedented reasoning capabilities; (2) neural scene representation frameworks [11] enabling photorealistic environment reconstruction; and (3) expanding real-world robot deployment across industrial and service sectors. The paper’s organizational architecture, visualized in Fig. 1, proceeds as follows:
  • Section II details two widely accepted definitions of PR: the position-based and the overlap-based definitions. It then offers a more precise formulation about “effective PR” and highlights the key challenges involved.
  • Section III reviews existing representation approaches in PR, covering the core solutions prevalent in the field. Intuitively, PR extends beyond mere image-based approaches, encompassing a variety of solutions. At a low level, detailed in Section III-A, a “place” can be captured through sensors, such as cameras, LiDARs and Radars. It’s generally expected that identical locations will produce similar sensor data. On a more abstract and higher level, as discussed in Section III-B, a “place” may also be represented through compact data forms, like scene graphs, implicit embeddings, and Gaussian Splatting (GS) maps.
@@
  • Section VI reviews the leading datasets and benchmarks in the field of PR. It introduces a new perspective on property analysis that complements to primary metrics for quality assessment.
@@
II Formulation of Effective Place Recognition and Challenges
@@
Before exploring specific solutions in PR, it is crucial to address two basic questions: (1) What is effective PR? and (2) What are the primary challenges encountered in PR?
II-A What is Effective Place Recognition?
II-A1 Existing Definitions
Two principal paradigms dominate PR definitions: position-based and overlap-based. Fig. 3 explains their differences in place judgement with an example. Originating from O’Keefe’s discovery of hippocampal “place cells” [12], position-based PR [2] evaluates whether a robot revisits a geographic location (point or region) despite environmental or viewpoint changes. The core challenge lies in robustly associating observations with spatial proximity. Conversely, overlap-based PR [5] defines place equivalence as visual overlap in the sensor’s field-of-view (FoV), irrespective of geographic distance, aligning with image retrieval [13].
While overlap-based PR mirrors content-based image search, its utility in robotic SLAM and navigation remains ambiguous. For instance, Fig. 3 illustrates two images observing the same landmark (e.g., a building) from distinct viewpoints: despite visual overlap, inferring their relative positions is non-trivial. Position-based PR better supports tasks like global localization [8], which prioritize coarse pose estimation. However, neither paradigm universally addresses all PR challenges. Thus, we emphasize effective PR as the intersection of two criteria: (1) Geographic proximity: places share meaningful spatial adjacency, and (2) visual consistency: observations exhibit measurable scene overlap or descriptor similarity.
II-A2 Formulation
Effective PR assumes a valid database image for the query image capture at place $Q$ must satisfy:
  • Geometric constraint: The candidate place $P\in\mathcal{P}$ is within a threshold distance $\delta$ from $Q$:
    $d(P,Q)<\delta,$ (1)
    where $d(\cdot,\cdot)$ measures translational or rotational distance.
  • Visual constraint: Their global descriptors $\mathbf{g}_{P}$ and $\mathbf{g}_{Q}$ exhibit high similarity with $\epsilon$ as a similarity threshold:
    $\|\mathbf{g}_{P}-\mathbf{g}_{Q}\|<\epsilon.$ (2)
This formulation excludes ambiguous cases (e.g., places on opposite sides of a wall or distant viewpoints observing the same landmark). A PR method succeeds only if it retrieves places satisfying both conditions, ensuring robustness to perceptual aliasing while maintaining spatial relevance.
@@ II-A2 Formulation
Figure 3: In position-based PR, the focus is on identifying if a query image stays at the same location as a database image. For instance, in the provided images, position-based PR would recognize the Query $1$ image as matching the database image, but reject the Query $2$ that is taken from a geographically distant location and considered as a different place. Overlap-based PR, however, would classify both query images as the same place since they share visual overlap with the database, denoted by the red box. But the Query $2$ offers limited utility for downstream navigation tasks.
@@ II-B Challenges
    Viewpoint Difference: This issue arises from the variations in how the environment is captured by sensors, which can be influenced by the sensor’s viewpoint, position, and intrinsic properties. For example, a building’s frontal view exposes its full shape, whereas a top-down perspective highlights the layout and roof design. Such viewpoint variations are especially noticeable when a robot revisits a location from a different angle or altitude. This challenge is ubiquitous across sensor types, necessitating PR systems to incorporate robust modules for feature extraction and matching to accommodate these perspective shifts.
@@
    Generalization Ability: For lifelong navigation, the vast complexity of environments makes generalizing to unseen areas a crucial challenge. PR methods must enable online learning to adapt over time, an essential requirement for missions like space exploration, where robots conduct long-term investigations on other planets with minimal human supervision. Robots need the ability to continuously learn and adjust to new environments.
@@
    Efficiency on Resource-Constrained Platforms: Deploying PR algorithms, particularly those requiring online fine-tuning, on resource-constrained robotic platforms is a major challenge, especially for aerial robots. Effective algorithms must strike a balance between accuracy and computational efficiency. This is essential for both single and multi-robot systems, which often encounter bandwidth limitations and potential communication disruptions.
@@
Building from above formulation and challenges, the following sections will delve into associate studies.
III Describing Places: Representation
The foundation of PR lies in how a place is represented as a map and then compared with incoming sensor data. Based on existing representation formats, we categorize these into low-level, sensor-specific representations and high-level, sensor-agnostic representations.
III-A Low-Level Representations
This approach represents a place as a database of primitive representations: images, point clouds, or features extracted from raw data with methods tailored to a specific sensor.
@@ III-A1 Sensor Selection Criteria
Cameras, LiDAR (Light Detection and Ranging), and Radar (Radio Detection and Ranging) are typical sensors in PR, as illustrated in Fig. 5. Important selection criteria include FoV, information density, and robustness under various conditions. Frame cameras provide high-resolution images, while event cameras perform better in low-light and reduce motion blur. LiDARs generate precise 3D point clouds but have low resolution. Radars offer long-range capabilities and excel in poor weather, also measuring relative velocity via the Doppler effect. Choosing the proper sensor or sensor combinations for PR depends on specific requirements for precision, range, and environmental suitability.
@@ III-A2 Camera-Related Approaches
VPR is the most thoroughly investigated problem. As highlighted in previous surveys (Section I-B), it has been comprehensively explored through both handcrafted and data-driven methodologies.
Handcrafted VPR Methods: Handcrafted representations in VPR are generally divided into local and global feature descriptors. Local descriptors such as SURF [22] identify a set of keypoints and compute corresponding descriptors for local image regions, but requiring extensive matching to compare visual similarity between images. In contrast, global descriptors, such as Gist [23], CoHOG [24], and BoW [25], aggregate local features into a unified descriptor, such as a vector or matrix, to capture an image’s overall pattern without local matching. More extensive review on tranditional methods can be referred to this survey paper [2]. While effective, these handcrafted methods generally underperform compared to data-driven, deep-learning approaches, which offer superior accuracy and robustness.
Data-Driven VPR Methods: Data-driven approaches [26, 27, 28, 29, 30, 31, 32, 33], particularly those utilizing deep neural networks (DNNs), automatically learn features from training data, reducing the need for extensive handcrafted design and domain expertise. A common two-step pipeline is often employed to enhance PR robustness: (1) initial matching and candidate selection by comparing global descriptors between the query and database data, and (2) re-ranking by refining matches using local features among selected candidates.
The emergence of CNN provided a new way to generate visual representations which had been proven to be successful in various category-level recognition tasks [34, 35]. The introduction of NetVLAD [26] represented a major step forward in data-driven VPR, using CNNs to transform images into feature maps, with a differentiable VLAD (Vector of Locally Aggregated Descriptors) pooling layer to create global descriptors. Later methods refined both feature extraction and aggregation, such as Regional Maximum Activations of Convolutions (R-MAC) [36] and Generalized Mean (GeM) [29, 30] as effective VLAD alternatives. Recent efforts also explored connecting global descriptors with local features.
Rather than focusing solely on enhancing local feature extraction from images, researchers are increasingly emphasizing the relationships between these features, which encapsulate valuable semantic information. Attention mechanisms and transformer architectures address this need by dynamically weighting feature relationships across the entire image, overcoming the constrained receptive field of CNNs while fostering more robust contextual semantics. Retriever [37], using cross-attention, and TransVPR [31] with multi-scale self-attention. Beyond transformers, MixVPR [28], inspired by isotropic all-MLP architectures like MLP-Mixer [38], proposes a novel all-MLP aggregation that performs competitively in one stage. SALAD [39] addresses the assignment in NetVLAD with an optimal transport approach using the Sinkhorn Algorithm [40].
Foundation models pre-trained on extensive datasets, like DINOv2 [41], exhibit strong zero- and few-shot generalizability for VPR. Methods such as AnyLoc [32] and SelaVPR [33] leverage DINOv2 to achieve state-of-the-art (SoTA) results. Additionally, PR results generated by global descriptors are still not accurate enough, especially for robot navigation tasks that require high top-$1$ recall. To enhance accuracy, several methods integrate local features. Patch-NetVLAD [27] pioneered re-ranking by generating sub-global descriptors using patches with NetVLAD. R2former [42] provides an end-to-end approach for computing re-ranking scores.
@@
Various research efforts have broadened scope by integrating diverse cues such as semantic, geometric, event data to boost the place representation. The utilization of semantics includes the way of filtering specific pixels [43] and adjusting the weight of feature embeddings [44]. Geometric cues, such as the 3D positions of landmarks, provide complementary structural information to visual descriptors and enhance PR accuracy [45]. Event cameras, with a higher dynamic range than frame cameras, have been also used in VPR to capture texture information in low-light conditions. Lee et al. [46] proposed to use event cameras to capture texture information under low-light condition, constructed edge-based images from event data to achieve PR.